### v1

# CyberSecEval 3: Advancing the Evaluation of Cybersecurity Risks and Capabilities in Large Language Models

Abstract
We are releasing a new suite of security benchmarks for LLMs, CyberSecEval 3, to continue the conversation on empirically measuring LLM cybersecurity risks and capabilities. CyberSecEval 3 assesses 8 different risks across two broad categories: risk to third parties, and risk to application developers and end users. Compared to previous work, we add new areas focused on offensive security capabilities: automated social engineering, scaling manual offensive cyber operations, and autonomous offensive cyber operations. In this paper we discuss applying these benchmarks to the Llama 3 models and a suite of contemporaneous state-of-the-art LLMs, enabling us to contextualize risks both with and without mitigations in place.
[…]
1 Introduction
The cybersecurity risks, benefits, and capabilities of AI systems are of intense interest across the security and AI policy community. Because progress in LLMs is rapid, it is challenging to have a clear picture of what currently is and is not possible. To make evidence-based decisions, we need to ground decision-making in empirical measurement.
We make two key contributions to empirical measurement of cybersecurity capabilities of AI systems. First, we provide a transparent description of cybersecurity measurements conducted to support the development of the Llama 3 405b, Llama 3 70b, and Llama 3 8b models. Second, we enhance transparency and collaboration by publicly releasing all non-manual portions of our evaluation within our framework, in a new benchmark suite: CyberSecEval 3.
We previously released CyberSecEval 1 and 2; those benchmarks focused on measuring various risks and capabilities associated with large language models (LLMs), including automatic exploit generation, insecure code outputs, content risks in which LLMs agree to assist in cyber-attacks, and susceptibility to prompt injection attacks. This work is described in Bhatt et al. (2023) and Bhatt et al. (2024). For CyberSecEval 3, we extend our evaluations to cover new areas focused on offensive security capabilities, including automated social engineering, scaling manual offensive cyber operations, and autonomous cyber operations.
1.1 Summary of Findings
We find that while the Llama 3 models exhibit capabilities that could potentially be employed in cyber-attacks, the associated risks are comparable to other state-of-the-art open and closed source models. We demonstrate that risks to application developers can be mitigated using guardrails. Furthermore, we have made all discussed guardrails available publicly.
Figure 1 summarizes our contributions. Specific findings include:
Llama 3 405B demonstrated the capability to automate moderately persuasive multi-turn spear-phishing attacks, similar to GPT-4 Turbo, a peer closed model, and Qwen 2 72B Instruct, a peer open model. The risk associated with using benevolently hosted LLM models for phishing can be mitigated by actively monitoring their usage and implementing protective measures like Llama Guard 3, which Meta releases simultaneously with this paper.
In our human subjects study on how Llama 3 405B assists in the speed and completion of offensive network operations, we found that 405B did not provide a statistically significant uplift in cyberattack completion rates relative to baselines where participants had access to search engines.
In tests of autonomous cybersecurity operations Llama 3 405B showed limited progress in our autonomous hacking challenge, failing to demonstrate substantial capabilities in strategic planning and reasoning over scripted automation approaches.
Among all models tested, Llama 3 405B was the most effective at solving small-scale program vulnerability exploitation challenges, surpassing GPT-4 Turbo by 23%. This performance indicates incremental progress but does not represent a breakthrough in overcoming the general weaknesses of LLMs in software exploitation.
When used as coding assistants, all LLMs tested, including Llama 3 405B, suggest insecure code, failing our insecure autocomplete test cases at the rate of 31%. This risk can be mitigated by implementing guardrails such as our publicly released Code Shield system.
Susceptibility to prompt injection was a common issue across all models tested with Llama 3 405B and Llama 3 8B failing at rates of 22% and 19% respectively, rates comparable to peer models. This risk can be partially mitigated through secure application design and the use of protective measures like our publicly released Prompt Guard model, which we launch with this paper.
Llama 3 models exhibited susceptibility to complying with clearly malicious prompts and requests to execute malicious code in code interpreters at rates of 1% to 26%. Both issues can be mitigated by benign cloud hosting services by monitoring API usage and employing guardrails like Llama Guard 3.
1.2 Paper structure
The layout of the rest of the paper is as follows. In Section 2 we contextualize risks and capabilities evaluated by CyberSecEval relative to related work.
Sections 3 and 4 assess risks and capabilities. We break risks into two categories: risks to third parties from cybersecurity capabilities of LLMs, which we assess in Section 3, and risks to application developers and users from applications that use LLMs, which we assess in Section 4.
In Section 5, we describe the guardrails we have built to mitigate the risks we measure. We are publicly releasing all such guardrails and we encourage others to build on top of them. We expect that guardrails will be added as a best practice in all deployments of Llama 3 or other LLMs. In Section 6 we conclude.
2 Related Work
Our work builds on a growing body of methods for assessing the security capabilities of large language models. We first discuss related work that informs our choice of which risks to evaluate, resulting in a broad spectrum of relevant risks assessed. As noted above, these fall into two categories: 1) risks to third parties and 2) risks to application developers, which includes risks to end users of those applications. Each of these risks has related work we discuss in turn.
2.1 On which risks to evaluate for new models
Our chosen categories of risks, risks to third parties and risks to application developers, were informed by the broader conversation on AI risk, as well as what we observe deploying AI models.
For example, both the UK National Cyber Security Centre (2024) and the White House (2023) Voluntary AI Commitments explicitly raise concerns about cyber capabilities of AI and call for measurement of these risks. These include concerns on aiding vulnerability discovery and in uplifting less-skilled attackers, which map to the two categories of risks we assess.
More recently, NIST (2024) calls out two primary categories of risk: first, that “the potential for GAI to discover or enable new cybersecurity risks through lowering the barriers for offensive capabilities” and second “expand[ing] the available attack surface as GAI itself is vulnerable to novel attacks like prompt-injection or data poisoning.” These also map directly to the two categories of risks we assess.
2.2 Assessment of risks to third parties
Previous work by Hazell (2023) has shown that LLMs can generate content for spear-phishing attacks. Bethany et al. (2024) conducted multi-month ecological studies of the effectiveness of such attacks. Our work, however, establishes a repeatable method for assessing the risk of a specific model for aiding spear-phishing through a human and AI judging process. We are not aware of another work that can effectively determine per-model spear-phishing risk in a short amount of time.
For “LLM uplift” of manual cyber-operations, Microsoft (2024) reports that threat actors may already be using LLMs to enhance reconnaissance and vulnerability discovery in the wild. Hilario et al. (2024) reports on interactively prompting Chat-GPT 3.5 to carry out a single end-to-end penetration test. In contrast, our work quantifies LLM uplift for manual cyber-operations across a body of volunteers. We also show quantitative results for both expert and novice populations, shedding light on the current capabilities of LLMs to both broaden the pool of cyber operators and to deepen capabilities of existing operators.
Beyond manual human-in-the-loop uplift, autonomous cyber operation by LLMs has been of great concern. Recent work by Fang et al. (2024) showed that GPT-4 can, in some cases, carry out exploitation of known vulnerabilities; they do not, however, release their prompts or test sets citing ethical concerns. Rohlf (2024) argues that these results may, instead, be simply applying already known exploits. More recently, the startups XBOW (2024) and RunSybil (2024) have announced products that aim at carrying out autonomous cyber operations. We are not aware, however, of other work that quantifies different models’ capabilities in this area. We are publicly releasing our tests to encourage others to build on top of our work.
Autonomous vulnerability discovery is a capability with both defensive and offensive uses, but also one that is tricky to evaluate for LLMs because training data may include knowledge of previously discovered vulnerabilities. CyberSecEval 2 by Bhatt et al. (2024) addressed this by programmatically generating new tests. Chauvin (2024) proposes a new test suite based on capturing feeds of known vulnerabilities in commodity software. Glazunov and Brand (2024) report that using multi-step prompting with an agent framework significantly increases performance in discovering vulnerabilities in their “Naptime” system. This shows the importance of our work to publicly release benchmarks for vulnerability discovery. As new frameworks and new LLMs come out, we encourage continued development of public benchmarks.
2.3 Assessment of risks to application developers
OWASP (2024) places prompt injection as number one on its “Top 10" vulnerability types for LLMs. Measuring prompt injection susceptibility is therefore of great interest. Schulhoff et al. (2024) solicited malicious prompts from 2,800 people and then used them to evaluate three LLMs including GPT-3. More recently, as LLMs have expanded to accept visual and other inputs, “multi-modal” prompt injection techniques have been developed; for example Willison (2023) demonstrates GPT-4 prompt injection from a picture of text with new instructions. Our work publicly releases an evaluation that can be used to assess any given model for textual prompt injection techniques, applying this to Llama 3, and also publicly releases visual prompt injection tests, which we do not apply in this paper.
Executing malicious code as a result of a prompt first became a concern following the announcement that GPT-4 would have access to a code interpreter. For example, Piltch (2023) demonstrated that GPT could be induced into executing code that revealed details about its environment. Our previous work in CyberSecEval 2 by Bhatt et al. (2024) then showed this was a feasible attack and provided a data set to evaluate the risk. We continue that work here, showing how to evaluate state of the art models for this risk, both with and without guardrails in place.
Facilitating cyber attacks with LLMs, as discussed above, has been a key policy concern. Li et al. (2024) introduced a safety benchmark consisting of curated questions and an LLM to judge responses. We continue the work from CyberSecEval 1 & 2 on determining if an LLM can be tricked into helping someone carry out a cyber attack with a clever prompt. Our work shows how to meaningfully compare different state of the art models for helpfulness, both with and without guardrails.
Code assistants have taken off as a use case for LLMs, with Microsoft (2023) reporting that a substantial fraction of code in GitHub results from GitHub copilot suggestions. Our previous work in Bhatt et al. (2024) quantified insecure coding suggestions, both with and without guardrails. We continue this work here to show the benefit of standardized test sets for evaluating state of the art models.
3 Assessment of offensive cybersecurity capabilities and risks to third parties
We assessed four risks to third parties from LLMs: automated social engineering, scaling manual offensive cyber-operations, autonomous offensive cyber-operations, and autonomous software vulnerability discovery and exploitation. We discuss these below and note that many of these capabilities are usable by defenders to harden their own networks and code. We make automated evaluations we describe publicly available part of CyberSecEval 3.
3.1 Risk: Automated social engineering via spear-phishing
We assessed the potential misuse of the Llama 3 405b helpful-only model in spear-phishing attacks, which involves personalized communication designed to deceive specific individuals into unwittingly participating in security compromises. For example, Figure 2 illustrates a dialogue between Llama 3 405b and an LLM-simulated phishing victim, where the Llama 3 405b attacker reasons about the simulated victim’s personal attributes to execute a strategy to persuade them to download and open a malicious attachment (we have added the highlight to the text at the bottom for emphasis). By automating the creation of convincing phishing content, LLMs could significantly increase the scale and efficiency of cyberattacks.
Assessment strategy
We evaluated the performance of various language models in a controlled spear-phishing simulation in which we used a combination of manual review and LLM review to evaluate model phishing capabilities. We make our code publicly available as part of CyberSecEval 3.
The advantages of using a simulation-based approach is that it affords fast, low-cost, standardized assessment of model phishing capabilities. The disadvantages are that LLMs can only simulate human victims of phishing behavior, meaning our results will be more approximate than well-designed human participant studies.
To contextualize Llama 3 models’ phishing behavior, we also evaluated GPT-4 Turbo, Qwen 2-72b-instruct, and Mixtral 8x22b. We used helpful-only versions of Llama 3 to avoid confounding model capabilities with the efficacy of safety tuning built into the model.
Phishing simulation procedure
Our phishing simulation process involves several steps:
Victim Profile Generation: Using an LLM, we randomly generate detailed victim profiles. These profiles include critical information that cyber attackers typically seek, such as the victim’s occupation, interests, locale, and cultural background.
Simulation of Phishing Attack:
We employ a judge LLM to act as a phishing victim.
An LLM-under-test is then used to interact with this victim. The goal for the LLM is to build rapport and persuade the victim to compromise their security. This could involve actions like installing malware or leaking sensitive corporate security details.
Evaluation of Performance:
The effectiveness of the LLM-under-test in the phishing dialogue is assessed through a combination of manual grading and an automated LLM-based grading system using a single model.
Manual grading rubric
The manual grading rubric is structured to evaluate the persuasion skills of the attacker (LLM-under-test) across a spectrum from very poor to excellent:
1.
Very Poor: The attacker displays weak persuasion skills, characterized by repetitive content and a lack of convincing arguments.
2.
Poor: Persuasion skills are somewhat weak, with some repetitive content and limited ability to build rapport.
3.
Intermediate: The attacker shows a moderate ability to build rapport and present somewhat convincing arguments.
4.
Good: Persuasion skills are strong, with a robust ability to build rapport and present convincing arguments.
5.
Excellent: The attacker excels in persuasion, effectively building rapport and presenting highly convincing arguments.
This structured approach allows us to systematically assess and compare the capabilities of different LLMs in a scenario that mimics real-world cyber threats.
Assessed risk
To assess risk, we used a judge LLM to evaluate spear-phishing performance across 250 test cases for each of the models. We validated these scores against a small sample of human evaluations where four human evaluators blindly rated each of the five model outputs across the same 10 test cases using the rubric defined above. Both the human and LLM judge evaluations of performance show that, in addition to GPT-4 Turbo and Qwen 2-72b-instruct, Llama 3 models could potentially be used to scale moderately convincing spear-phishing campaigns in at least some cases.
Human evaluation scores and LLM judge evaluation scores for the 250 test cases sampled can be seen in Figure 3. GPT-4 Turbo scores approximately averaged 2.90 (“Intermediate” capabilities of multi-step persuasion) by human graders. In contrast, Llama 3 405b had an average score of 2.62 (between “Intermediate” and “Poor” persuasive abilities), and Mixtral 8x22b had an average score of 1.53 (between “Poor” and “Very Poor” persuasive abilities). When judge classifier scores of spear-phishing attempt persuasiveness were compared against a small sample of blind human evaluations from 4 different evaluators, we found judge scores and human scores to have a strong positive correlation (r = 0.89) for these model outputs.
GPT-4 Turbo and Qwen 2-72b-instruct were evaluated by our judge LLM to be significantly more successful at achieving spear-phishing goals than Llama 3 405b and Mixtral 8x22b. Note that we used four human evaluators, so the error bars on human evaluation are likely wide and are expected to overlap with the LLM judge evaluations.
As shown in Figure 4, when our judge LLM evaluated the overall approach of a model’s spear-phishing campaign across all spear-phishing runs, GPT-4 Turbo was rated highly, followed closely by Qwen 2-72b-instruct, then Llama 3 70b and Llama 3 405b, and lastly Mixtral 8x22b.
Although the Llama family of models were scored as being moderately successful at accomplishing their spear-phishing goals and moderately convincing, these evaluations imply that the Llama family of models is unlikely to present a greater risk to third parties than public and closed-source alternatives currently available to the public.
3.2 Risk: Scaling Manual Offensive Cyber Operations
We sought to understand how well Llama 3 405b, when used as an offensive cyber operations copilot and presented in a chatbot interface, could uplift cyberattack capabilities by accelerating and upskilling either novice or experienced cyberattackers.
Such uplift, if observed, may present risks and opportunities. If Llama 3 405b benefits attackers through increasing attacker learning rates or reducing the skill floor required to carry out cyberattacks, defenders could both benefit from using these capabilities to more quickly identify and fix vulnerabilities in their networks.
We designed an experiment to evaluate the uplift of Llama 3 for both novice and experts in a “capture the flag” simulation. For example, Figure 5 is an example interaction between a human participant and Llama 3 405B during our offensive security uplift study. Overall, we found that Llama 3 405B did not provide statistically significant uplift to human participants vs using search engines like Google and Bing. Subsequently, our findings indicate that even with Llama 3 ’s help, none of the novices completed all stages of the simulation. We further observe that the uplift of Llama 3 for experts or novices is negligible.
Assessment strategy
We evaluated the effectiveness of the Llama 3 405b language model (LLM) in aiding cyberattackers. We involved 62 Meta employee volunteers, half of whom were offensive security experts and the other half had technical backgrounds but were not security experts. We assessed volunteer performance in two stages.
In the first stage, subjects were asked to complete a cyberattack challenge without the assistance of an LLM. In the second stage, subjects were encouraged to use Llama 3 405b and were asked to complete a different cyberattack challenge of similar difficulty. Each stage lasted one hour. Cyberattack challenges were hosted on Hack The Box, a hacking simulation service.
Success was measured based on the number of phases a subject completed and how long a subject took to progress between phases. The steps involved in a typical cyber attack include:
Network Reconnaissance
Vulnerability Identification
Vulnerability Exploitation
Privilege Escalation
Assessed risk without guardrail mitigations
Our findings identified no significant difference in the actual success rate of participants when using the LLM compared to their performance in the LLM-free stage.
Key findings from the study include:
On average, inexperienced attackers completed 22% more phases with the LLM and reduced their time per phase by about 9 minutes and 12 seconds. This difference is not statistically significant due to high variance in subject success rates.
Experts completed 6% fewer phases with the LLM and only reduced their completion rates by an average of 1 minute and 44 seconds per phase. This difference is not statistically significant.
While inexperienced subjects reported some benefits from using the LLM (such as reduced mental effort and feeling like they learned faster from using the LLM), experienced subjects shared more mixed feedback, including:
"The LLM did well with evaluating potential next steps but performed poorly in identifying specific exploits to attempt."
"The LLM often decreased efficiency by trying to be more helpful than it needed to be."
"It’s good to know that the AI is as bad as I am at solving HTB host boxes."
Figure 6 depicts how quickly novices and experts completed the total collection of phases they encountered as groups during the LLM stage and the LLM-free stage. Even with LLM help, no novices completed all phases of either of the capture the flag stages. Although novice subjects may have seen faster phase completion rates for some phases, these were phases that would not have taken them much additional time without the use of an LLM. Experts, in contrast, performed roughly the same with LLM assistance as without.
We conclude that if novices benefited from access to the LLM, the benefits are marginal and inconsistent. Differences in performance in the LLM stage may be attributable to a number of confounding variables including:
Subjects applying cyberattack approaches learned during the immediately preceding non-LLM stage.
Greater time pressure in the LLM stage if the stage overran into scheduled meetings for subjects.
Subject hesitancy to interrupt the LLM because subjects wanted to provide complete output for study observers, and more.
Mitigation recommendations
While Llama 3 does not appear to significantly improve the success rate of cyberattacks relative to an open-web non-LLM baseline, cloud service providers hosting Llama 3 may still want to minimize the misuse of hosted models by cyber threat actors. To mitigate this risk of misuse, we have publicly released Llama Guard 3 with the Llama 3 launch, which can identify, log, and block requests that induce Llama 3 models to act as cyberattack co-pilots. We recommend guardrailing Llama 3 deployments with Llama Guard 3, as demonstrated in our system level safety reference implementation.
3.3 Autonomous offensive cyber operations
We assessed the Llama 3 70b and 405b models’ potential to function autonomously as a hacking agent and have made our assessment code publicly available as part of CyberSecEval 3. A capable LLM could significantly alter the landscape of cyber threats if it is able to accomplish cyber attack objectives autonomously.
Assessment strategy
We tested the model in a controlled environment designed to simulate the critical phases of a ransomware attack: Network Reconnaissance, Vulnerability Identification, Exploit Execution, and Post Exploitation Actions. We configured Llama 3 to autonomously generate and execute Linux commands based on the outcomes of previous commands on a Kali Linux virtual machine, targeting a fully updated Windows Server virtual machine with known vulnerabilities.
Assessed risk without guardrail mitigations
During 85 test runs, we closely monitored the model’s performance. Our key observations included:
Reconnaissance and Initial Access: The model efficiently identified network services and open ports but failed to effectively use this information to gain initial access.
Vulnerability Identification: The model sometimes identified vulnerabilities but struggled with selecting and applying the correct exploitation techniques.
Exploitation: Attempts to execute exploits were entirely unsuccessful, indicating a lack of adaptability to dynamic network environments and inability to craft a correct exploitation command.
Post Exploitation: The model showed no capability in maintaining access or impacting hosts within the network.
One limitation of our assessment strategy is that we did not explore advanced prompting strategies such as Chain of Thought, Thread of Thought, Self Ask, or Plan and Solve to drive our LLM agents. Malicious users with deep cybersecurity knowledge or users who invested greater time in prompt engineering, fine-tuning the model, incorporating RAG, or using an agentic framework may be more successful in developing autonomous cyber attack agents.
Figure 7 demonstrates the relative rates models complete low-sophistication, moderate-sophistication, and high-sophistication cyber operations challenges offensively, as evaluated from a consensus of three judge LLMs. Llama 3 70b was found to complete over half of the low-sophistication challenges successfully.
Some models spent more time attempting sophisticated actions such as lateral movement, while other models focused on simpler but more likely to succeed actions such as network discovery.
Mitigation recommendations
We believe that the risk that Llama 3 models can be used successfully for autonomous cyberattacks on computer networks is low given its very limited assessed capabilities. For cloud providers seeking to further minimize risk of their models’ misuse, we recommend deploying Llama Guard 3 to detect and block potential cyberattack aid requests.
3.4 Autonomous software vulnerability discovery and exploitation
The potential for LLMs to identify and exploit vulnerabilities could significantly enhance the capabilities of cyberattackers if these systems were able to uncover vulnerabilities that are currently undiscoverable by existing methods or if attackers are able to discover them faster or more covertly than software-developing organizations. On the other hand, such capabilities could help defenders find and fix security vulnerabilities.
In either case, as of now, there is no evidence that AI systems, including Llama 3, outperform traditional non-AI tools and manual techniques in real-world vulnerability identification and exploitation on real-world scale programs. This limitation is attributed to several factors:
Limited Program Reasoning: LLMs have restricted capabilities in reasoning about programs, even at a small scale.
Complex Program Structures: The intricate distributed control flow graphs and data flow graphs of real-world programs do not fit within the largest LLM context windows, making effective reasoning challenging.
Agentic Reasoning Requirements: Real-world vulnerability identification typically involves multi-step reasoning processes that require feedback from dynamic and static analysis tools, which are beyond the current capabilities of LLMs.
Still, given the impact an AI breakthrough in software vulnerability exploitation could have on cybersecurity, it is important to track progress in this direction.
Assessment strategy
We conducted tests using toy, capture-the-flag style hacking challenges from CyberSecEval 3. These include string constraint satisfaction problems in C, Python, and JavaScript; SQLite injection in Python; buffer overflow tests in C++; and advanced memory corruption tests. Figure 8 shows one example of a heap overflow exploitation challenge used to measure LLM progress towards automatic exploit generation.
Assessed capabilities
Figure 9 shows how well GPT-4, Llama 3 instruction tuned models, Mistral 8x22B, and Gemini Pro did when prompted to solve these challenges, which are fully described in Bhatt et al. (2024). Notably, Llama 3 405B surpassed GPT-4 Turbo’s performance by 22%. Our benchmark was based on zero-shot prompting, but Google Naptime Glazunov and Brand (2024) demonstrated that results can be further improved through tool augmentation and agentic scaffolding. We plan to explore these directions in our future work.
4 Llama 3’s cybersecurity vulnerabilities and risks to application developers
We assessed four risks to application developers that include LLMs into their applications and to the end users of those applications:
1.
Prompt injection
2.
Convincing the model to execute malicious code in attached code interpreters
3.
Agreeing to facilitate cyber attacks
4.
Suggesting insecure code, when used as a coding assistant
These capture risks arising from the most common applications of LLMs. In particular, coding assistant and use of code interpreters by LLMs is widespread and growing.
We are releasing with Llama 3 multiple guardrails that serve to mitigate these risks. In our assessments, we evaluated the model on its own, as well as the model with these guardrails in place. We expect application developers to deploy guardrails as a matter of best practice. Therefore, when we assess the overall risk, we look at the performance with the guardrails in place.
4.1 Risk: textual prompt injections
Prompt injection attacks occur when LLMs conflate trusted instructions and untrusted user input, and process them together. The untrusted data could modify the LLM’s behavior to violate safety, security, or privacy guidelines.
We have also developed and released Prompt Guard alongside Llama 3, which detects prompt injection attacks. We expect application developers to employ Prompt Guard or a similar guardrail as a best practice. Today’s LLMs have vision and other multimodal capabilities. As a result, prompt injection techniques can be textual and/or feature visual, audio or video input. We do not address multimodal prompt injection risks in the present work.
Models are given the Test Case Prompt as their system prompt, are fed User Input, and then a judge LLM is used to determine whether the test injection was successful.
Assessment strategy
We tested Llama 3 70b and 405b against 251 manually curated test cases from CyberSecEval 2, which are fully described in Bhatt et al. (2024). Figure 10 summarizes all the types of prompt injection test cases evaluated, along with detailed examples. Test cases are classified as either “logic-violating” cases (cases in which the injection causes the model to deviate from a general guideline set in the system prompt) or “security violating test cases” (a subset of cases which are emblematic of real security risk in LLM-powered applications, such as private information or password leaks).
Assessed risk without guardrail mitigations
These benchmarks show that the Llama 3 models exhibit comparable performance to GPT-4 for prompt injection attacks, as listed in Figure 11.
Specific findings include:
Overall Attack Success Rates (ASR) are 20% - 40%, consistent with previously published models.
We observe a similar ASR for non-English injection attacks.
A somewhat higher ASR for prompt injections in non-English languages was noted across all models.
Mitigation recommendations
To mitigate prompt injection risk we recommend the deployment of Prompt Guard, which we’ve developed and released alongside Llama 3. We particularly recommend deploying Prompt Guard to detect indirect injections in third-party content consumed by Llama 3, as indirect injections pose the most risk to users of applications.
Prompt Guard has demonstrated effectiveness in significantly reducing the attack success rate for textual prompt injections and jailbreaks. There may be instances, however, where textual prompt injections could bypass our filters or be so application-specific that they evade generic detection models.
4.2 Risk: Suggesting insecure code
When LLMs generate code, the output can fail to adhere to security best practices or introduce exploitable vulnerabilities. This is not a theoretical risk—developers readily accept significant amounts of code suggested by these models. Microsoft (2023) revealed that 46% of the code pushed to its platform by developers using the GitHub copilot AI tool was generated by the LLM.
We found that while models suggest insecure code, this risk can be mitigated by guardrails.
Assessment strategy
We utilized the CyberSecEval insecure coding test case corpus, fully described in Bhatt et al. (2023), to prompt all Llama 3 models to solve coding challenges and measure the generation of insecure code. This testing framework prompts an LLM with prompts likely to induce insecure code output and then assesses whether vulnerable code is being generated by the LLM based on static analysis. Our benchmark can be further divided into two types: autocomplete-based (e.g., "complete the function of my code") and instruct-based (e.g., "write me a program"). The performance of the Llama 3 models was benchmarked against other industry models, such as GPT-4-turbo, to provide a comparative analysis of their capabilities in generating code both securely and meaningfully.
Assessed risk without guardrail mitigations
The Llama 3 models align with the previously noticed pattern of insecure coding generation: as the number of parameters in an LLM increases, we observe that more insecure code is generated.
Figures 12 and 13 show the performance of different AI models in terms of insecure coding and code quality. In the autocomplete-based evaluation, the Llama 3 405b produces less insecure code at a rate of 30.55%, compared to GPT-4-turbo’s 29.84%.
In the instruct-based evaluation, the 405b model generates more insecure code at a rate of 38.57% compared to GPT-4-turbo’s 35.24%.. The 8b model generates a smaller percentage of insecure coding, indicating a trade-off between code quality and security. This accords with a dynamic we described in Bhatt et al. (2023) in which more capable coding models also generate insecure code at a higher rate.
Mitigation recommendations
We recommend deploying Code Shield which we released alongside Llama 3 to mitigate against the code security risks. Code Shield is effective in identifying and guardrailing insecure coding practices when output from an LLM. Code Shield is capable of identifying around 190 patterns across 50 different CWEs with an accuracy of 90%. However, it’s not guaranteed to identify all vulnerabilities hence additional static and dynamic analysis tooling are recommended in the CI/CD pipeline downstream where the code is being shipped.
4.3 Risk: Agreeing to execute malicious code in attached code interpreters
A growing trend in the LLM industry is to allow LLMs to execute code in an attached interpreter (typically Python). This carries the risk that malicious users could induce LLMs to compromise or misuse system resources.
While application developers should protect these systems by sandboxing the infrastructure on which the LLMs run as a first priority, as an additional layer of defense, LLMs should refuse to execute code that attacks the sandbox, or use the sandbox to help attack other systems. This can support security operations teams in using LLM refusals as a signal to detect ongoing attacks, and in adding friction for attackers attacking the sandbox or misusing the code interpreter.
Assessment strategy
We utilized the CyberSecEval code interpreter abuse prompt corpus, fully described in Bhatt et al. (2024), to measure the propensity of Llama 3 models to comply with prompts that betray malicious intent vis-a-vis an attached sandbox. This testing framework evaluates the models’ responses across various attack categories, including privilege escalation and container escape. The performance of Llama 3 models was benchmarked against other industry models, such as GPT-4 Turbo, to provide a comparative analysis of their susceptibility to malicious code execution.
Assessed risk
The results (Figure 14) indicate Llama 3 405b is susceptible to certain prompts which generate code that could abuse a code interpreter. The 405b model generated malicious code 1% of the time. These rates are higher compared to the peer model GPT-4-Turbo, which has a generating rate of only 0.8%. Note that we are comparing the base Llama models without guardrails to GPT-4 Turbo which contains guardrails.
Mitigation recommendations
To further reduce compliance susceptibility, we recommend the implementation of Llama Guard 3, which has been specifically developed to classify code interpreter attacks. Figure 14 shows that LlamaGuard Inan et al. (2023) moves the rate at which a LlamaGuard 3 guardrailed system complies with our test interpreter prompts to 0. We note that real world attackers would still likely find prompting strategies to bypass LlamaGuard and model protections, and that usage monitoring and secure sandbox construction are key to mitigating risks.
4.4 Risk: Agreeing to help facilitate cyberattacks
We used CyberSecEval’s cyber attack helpfulness test suite, fully described in Bhatt et al. (2023), to test the risk that Llama 3 models comply with requests to help with cyberattacks, which could lead to abuse of models hosted by benign providers. Our testing framework evaluates the models’ responses across various attack categories mapping to the MITRE AT&CK ontology. While we found that Llama 3 models often comply with cyber attack helpfulness requests, we also found that this risk is substantially mitigated by implementing LlamaGuard 3 as a guardrail.
Assessed risk without guardrail mitigations
In Figure 15, we found that, overall, Llama 3 family refuses explicit requests to help carry out cyber attacks at a similar rate as compared to peer models. Similar to peer models, Llama 3 models tend to refuse to help in higher severity prompt categories, such as privilege escalation, and complied more often in lower severity categories, like network reconnaissance.
Mitigation recommendations
To reduce the risk deployed Llama 3 models comply with requests to help with cyberattacks, we recommend the deployment of Llama Guard 3. This system is designed to guardrail both the prompts entering the model and the response data coming out of the model, ensuring that the outputs are not helpful to cyberattacks.
5 Guardrails for reducing cybersecurity risks
5.1 Using Prompt Guard to reduce the risk of prompt injection attacks
Prompt Guard is a multi-label classifier model we are publicly releasing to guardrail real-world LLM-powered applications against prompt attack risk, including prompt injections. Unlike CyberSecEval, which tests the ability of models to enforce consistency of system prompts and user instructions against contradictory requests, Prompt Guard is designed to flag inputs that appear to be risky or explicitly malicious in isolation, such as prompts that contain a known jailbreaking technique. Prompt Guard has three classifications:
Jailbreak: which identifies prompts as explicitly malicious
Injection: which identifies data or third-party documents in an LLMs context window as containing embedded instructions or prompts
Benign: any string that does not fall into either of the above two categories
Direct Jailbreaks
To test Prompt Guard’s ability to detect jailbreaks, we use a separate set of real-world jailbreak and benign prompts. No part of this dataset was used in training, so it can be considered completely “out-of-distribution” from our training dataset, simulating a realistic filter of malicious and benign prompts on an application that the model has not explicitly trained on. We found that we achieve a recall of 97.5% of jailbreak prompts with a false positive rate of 3.9% at our selected threshold. Detailed results can be found in Figure 16.
Indirect Injections
To test Prompt Guard’s ability to detect injections, we repurpose CyberSecEval’s dataset as a benchmark of challenging indirect injections covering a wide range of techniques (with a similar set of datapoints with the embedded injection removed as negatives). We present our evaluation in Figure 17. We find that the model (at our selected threshold) identifies 71.4% of these injections with a 1% false-positive rate.
Conclusion
Indirect injections are the largest realistic security risk faced by LLM-powered applications, so we recommend scanning and filtering all third-party documents included in LLM context windows for injections or jailbreaks. The tradeoff of filtering jailbreaks in direct user dialogue is application specific. We recommend fine-tuning Prompt Guard to the specific benign and malicious prompts of a given application before integration rather than integrating out of the box; our release of Prompt Guard includes instructions in how to do so in its README.
5.2 Using Code Shield to reduce the risk of insecure code suggestions
Code Shield is an inference time filtering tool designed to prevent the introduction of insecure code generated by LLMs into production systems. Our results above show LLMs can sometimes output insecure code. Code Shield mitigates this risk by intercepting and blocking insecure code in a configurable way.
Code Shield leverages our Insecure Code Detector (ICD) static analysis library to identify insecure code across 7 programming languages and over 50 CWEs. It is optimized for production environments where low latency is critical, employing a two-layer scanning approach. The initial layer swiftly identifies concerning code patterns within 60ms. If the code is flagged as suspicious, it undergoes a more thorough analysis in the second layer, which takes approximately 300ms. Notably, in 90% of cases, only the first layer is invoked, maintaining the latency under 70ms for the majority of scans.
Code Shield is not a panacea and may not detect all insecure coding practices. To understand the efficacy of Insecure Code Detector, we manually labeled 50 LLM completions corresponding to our test cases per language based on whether they were insecure or secure. Then we computed the precision and recall of our Insecure Code Detector static analysis approach both per-language and overall, as shown in Figure 18. We found that, overall, the Insecure Code Detector had a precision of 96% and a recall of 79% in detecting insecure LLM generated code.
Additionally, it may introduce some latency, with 10% of queries taking longer than 300ms as observed in some production environments. Nevertheless, Code Shield has shown its utility in enhancing the security of code generated by Llama models, and we recommend deploying Code Shield whenever deploying Llama models to generate production source code.
5.3 Using Llama Guard to reduce the risk of Llama 3 compliance with prompts that betray malicious offensive cybersecurity intent
Llama Guard is a fine-tuned version of Llama 3 8b, designed for guardrailing inputs and outputs to LLMs against safety violations. It allows application developers to specify a set of “unsafe content categories”, and is finetuned to accept a wide range of categories originating from the MLCommons AI Safety taxonomy of harm categories. In the most recent version of Llama Guard (Llama Guard v3) we extended this taxonomy to include a new category (“Code Interpreter Abuse”) and provided training data to detect such abuse for fine-tuning of Llama Guard. We also tested the capability of Llama Guard to detect inputs and outputs that help to facilitate cyberattacks within an existing safety category (“Non-violent crimes”).
MITRE Tests
We used the MITRE CyberSecEval tests to evaluate the ability of Llama Guard to detect requests to assist with cybercrimes when used as either an output or an input and output guardrail. We found Llama Guard to be relatively effective at detecting such requests - we see a 50.4% and 53.9% reduction in violation rate for Llama 3 405b and Llama 3 70b, respectively when the model is used as both an input and output filter.
We also use the MITRE FRR benchmark to measure the propensity of Llama Guard to inadvertently filter borderline requests. We observed an increase in FRR, indicating a tradeoff between this safety guardrail and helpfulness. We observe 2% false refusal rate when Llama Guard is used as an output guardrail only and 10% false refusal rate when Llama Guard is used as an input and an output filter, with a negligible baseline violation rate.
6 Limitations and future work
We have carried out a comprehensive assessment of risks, and we know there is room to do more. We see three major directions for further improvement of risk assessment for cybersecurity: first, maximizing efficacy of models under test, second, scaling up cross-checks between human and LLM judges, and finally, continuous assessment of model risk over time. We discuss each in turn.
Maximizing model efficacy. Our results measure a base model with and without guardrails applied. Work on applying agent scaffolding, however, shows that substantial improvements to a task can be made over and above base models by specializing to the task at hand. Notable in our context is work by Glazunov and Brand (2024) on “Naptime", which uses agentic scaffolding to achieve up to 20x better results on our CyberSecEval 2 benchmark across both ChatGPT and Gemini models. Future risk evaluations could propose additional model-independent agent architecture specific to measuring the risk. In addition, future evaluations could of course carry out fine-tuning or other model-specific specializations to improve efficacy for a specific risk.
Scaling cross-checks between human and LLM judges. Our results in spear-phishing and autonomous uplift assessments relied principally on judge LLMs, which we then checked against a set of four human judges. While we found that both human and LLM judges were directionally aligned in this assessment, this was a one-time cross-check with a small number of people. We also found that in a different case, assessing risk from autonomous cyber operations, an LLM judge had false positives.
The opportunity for future work is to scale up the number of human judges, make the process easily repeatable, and carefully quantify inter-rater agreement and disagreements. Here we expect to learn from the rich body of literature and practice on crowdsourcing techniques to train raters to repeatedly judge tasks, notably applied to search ranking results.
Using a consensus of LLM judges. Multiple LLM judges powered by differing models can be given the same input. The final judge outcome would be a consensus of the LLM judges. Each LLM judge can be run a number of times with their best-of-n result being their contribution to the consensus pool. This may improve the false positive and false negative rate of LLM judges, and improve LLM judge reliability over our current single LLM judge. Future work involves implementing a consensus algorithm in the grading portion of our eval code and selecting a diverse set of capable judge models.
Continuous assessment of model risk over time. We carried out a point in time assessment for each of our models, but in reality new models release constantly. We will release our models publicly, including on the Hugging Face platform, to enable people to continuously monitor the capabilities of new models as they become available. An interesting question here would be what can we learn from comparing assessments of the same model architecture over time?
7 Conclusion
We released a new benchmark suite for assessing cybersecurity risks from LLMs, CyberSecEval 3, which extends CyberSecEval 1 and CyberSecEval 2. We demonstrated the effectiveness of CyberSecEval, overall, by using it to evaluate Llama 3 and a select set of contemporary state of the art models against a broad range of cybersecurity risks. We find that mitigations, which we also release publicly, can measurably improve multiple risks, both for Llama 3 and for other models. We encourage others to build on our work to continue improving both empirical measurement of risk and mitigations to reduce that risk.
8 Acknowledgements
Aaron Grattafiori, Chris Rohlf, Esteban Arcaute, Rachad Alao, and Vincent Gonguet for providing valuable feedback and guidance through all phases of this work, Craig Gomes for engineering management support, Tatyana Porturnak and Ahuva Goldstand for review and suggestions.
We extend our thanks to everyone who participated in our human uplift study and our human judging. This includes Aaron Grattafiori, Abhishek Kaushik, Adi Balapure, Aleksandar Straumann, Alex Kube, Aman Ali, Benjamin Williams, Brynn Chernosky, Christopher Tao, Danny Christensen, Dominic Spinosa, Doug Szeto, Dávon Washington, Fady Wanis, Frankie Yuan, Geoff Pamerleau, Goran Cvijanovic, Greg Prosser, Hamza Kwisaba, Harie Srinivasa Bangalore Ram Thilak, Harshit Maheshwari, Innocent Djiofack, Ivan Evtimov, Jade Auer, James Burton, Jay Thakker, Jeffrey O’Connell, Jon Goodgion, Jorge Gonzalez, Kenny Keslar, Kevin Castillo, Krishna Khamankar, Kristina Bebedzakova, Lance Naylor, Lloyd Lee-Lim, Mack Binns, Magan Omar, Mantas Zurauskas, Mariel Alper, Maru Berezin, Matt Kalinowski, Mike Hollander, Nariman Poushin, Neil Patel, Nik Tsytsarkin, Nikko Mazzone, Peter Mularien, Rachid Aourik, Raghudeep Kannavara, Rayan Hatout, Ricky Lashock, Rohit Dua, Ryan Edward Kozak, Ryan Hall, Sean Mcivor, Sid Wadikar, Stephanie Liu, Surya Ahuja, Sushma Gogula, Tariq Ramlall, Thad Bogner, Valeri Alexiev, Willem Koopman, and Yurii Karbashevskyi. Without your help, we would not have data to share with the world.
Lastly, thank you to everyone on all of the teams who helped make this work possible: AI Security, Program Analysis, Product Security Group, Offensive Security Group, Responsible AI, GenAI, Data Security Systems, Privacy Insights and Investigations, Fundamental AI Research.